• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Holmstrom

    Log entry time 18:04:14 on July27,2003

    Entry number 107483

    keyword=Kinimatic Switch

    We are done with kinimatic 6.16, and we are switching to kinimatic 6.18.
    We have set the Septum set current to 159 A, and the right HRS magnets to 1.707.